# Clarify Confused Nodes via Separated Learning

### v3

@@ VI-A3 Setup
To ensure the stability and reproducibility of the results, we use ten consecutive seeds to fix the data splitting and model initialization. For all methods, we set the maximum epochs to 500 with 100 epochs patience for early stopping, the hidden dimension to 512, and the optimizer to Adam [50]. For common parameters like learning rate, weight decay and dropout rate, we search in the same parameter space if their codes are publicly available. The key parameters of NCGCN/NCSAGE like hop and threshold will be searched in their respective parameter spaces and . Detailed parameter settings are reported in Appendix -C.
@@ VI-B Analysis
Table I reports the node classification results of all methods on ten benchmarks, from which we can draw the following conclusions: 1) Among all baselines, spectral methods show a dominant performance advantage on the heterophily problem. However, although they achieve relatively high overall performance, they lack a certain degree of universality, e.g., GPRGNN does not perform well on Photo and Computer. This is mainly due to the fact that the high-order prior of these methods is not applicable to all datasets, and the weight sharing hinders further performance improvement. 2) Spatial methods usually have complex architectural designs and large hyperparameter spaces, leading to higher computational complexity as well as unstable and poor performance. For instance, WRGAT, GBKSage and HOG are out of memory (OOM) on Coauthor Physics and Penn94. 3) As a spectral method, our framework achieves the best overall performance on two types of datasets without being out of memory (OOM), and reaches first place on the public leaderboard in six well-explored benchmarks (refer to Appendix -D), indicating its performance advantage. Additionally, our framework shows an average performance improvement of 8.04% and 4.54% on their respective backbones (GCN and SAGE), confirming its general plug-in property. Notably, the sparsification of the adjacency matrix by the NC-mask allows our NCSAGE model to overcome the OOM problem encountered by SAGE on Penn94. All these phenomena highlight the effectiveness of our framework in alleviating heterophily problem, especially the superiority of our proposed NC metric and priors.
@@ VI-B5 RQ5: How do the two hyperparameters of hops and threshold impact NCGCN?
Metric computation and node separation in our framework are controlled by the two key hyperparameters, hops and threshold . To further investigate the impact of these hyperparameters on NCGCN, we fix the other parameters to their optimal settings and plot the test accuracy curves under varying and , as shown in Fig. 9. It can be seen that NCGCN is less sensitive to hyperparameter changes on homophilous graphs than on heterophilous graphs, as the former usually contains relatively fewer high-NC nodes. On heterophilous graphs, NCGCN exhibits varied hyperparameter sensitivities across different datasets. First, on heterophilous graphs, using 1-hop information to measure the neighborhood confusion of nodes generally yields better overall performance and lower sensitivity than using 2-hop information. This is likely due to the more complex label distribution in the 2-hop neighborhood. Second, choosing different thresholds for node separation significantly impacts model performance, as nodes with varying degrees of NC values are often distributed in heterophilous graphs (see Fig. 5). For small-scale heterophilous graphs like Actor, Squirrel Filtered and Chameleon Filtered, NCGCN is more sensitive to threshold changes due to the more discrete distribution of nodes across different NC values. For larger heterophilous graphs like Penn94, the distribution of nodes with different NC values trends towards increasing numbers with higher NC values, leading the model to exhibit a monotonic performance fluctuation trend in response to threshold changes when using 1-hop information.
@@ VI-B7 Efficiency Comparison
First, CPGNN, GBKSage and HOG employ two-phase training and soft rewiring techniques, resulting in significant time consumption. WRGAT constructs a weighted multi-relational graph before training, rendering it equivalent to multi-channel GATs with reduced efficiency. These spatial methods typically involve additional computational steps for adjusting message passing rules, such as rewiring, auxiliary tasks, or gating mechanisms, thereby increasing the computational burden. In contrast, the NCGCN architecture is essentially a simple deformation of GCN, while also incorporates the three key designs mentioned in Sec. V-D. Consequently, it avoids excessive computational overhead, ultimately ensuring good computational efficiency. Please refer to Appendix -E for more analysis. Additionally, although some spatial methods consume less computation time, they have certain design flaws in their model architectures. For GloGNN, directly inputting the adjacency matrix into an MLP contradicts permutation invariance. For SNGNN, the approach of selecting the most similar neighbors seems to remain confined within the homophily assumption. CAGNN combines raw and aggregated features without considering label heterogeneity. These design flaws result in these methods falling short in terms of accuracy.
@@ VII Conclusions and Outlook
Our framework is simple yet effective enough to be a new start for spatial GNNs. The backbone can be replaced by more message passing models. The robustness and debiasing potential are also open to be explored. Furthermore, NC is an easy-to-compute metric that can effectively separate nodes for different optimization processes. This metric can also be further utilized in data mining, predictability, and interpretability. The limitation of NCGCN is that it utilizes only second-order information following classical GNNs. It seems sufficient on most real-world datasets except for long-distance dependent ones. Another problem is that high-NC nodes are relatively rare on homophilous graphs, and we can use graph data augmentation [51] for this issue in the future.

### v6

@@ VI-A3 Setup
To ensure the stability and reproducibility of the results, we use ten consecutive seeds to fix the data splitting and model initialization. For all methods, we set the maximum epochs to 500 with 100 epochs patience for early stopping, the hidden dimension to 512, and the optimizer to Adam [55]. For common parameters like learning rate, weight decay and dropout rate, we search in the same parameter space if their codes are publicly available. The key parameters of NCGCN/NCSAGE like hop and threshold will be searched in their respective parameter spaces and . Detailed parameter settings are reported in Appendix -D.
@@ VI-B Analysis
Table I reports the node classification results of all methods on ten benchmarks, from which we can draw the following conclusions: 1) Among all baselines, spectral methods show a dominant performance advantage on the heterophily problem. However, although they achieve relatively high overall performance, they lack a certain degree of universality, e.g., GPRGNN does not perform well on Photo and Computer. This is mainly due to the fact that the high-order prior of these methods is not applicable to all datasets, and the weight sharing hinders further performance improvement. 2) Spatial methods usually have complex architectural designs and large hyperparameter spaces, leading to higher computational complexity as well as unstable and poor performance. For instance, WRGAT, GBKSage and HOG are out of memory (OOM) on Coauthor Physics and Penn94. 3) As a spatial method, our framework achieves the best overall performance on two types of datasets without being out of memory (OOM), and reaches first place on the public leaderboard in six well-explored benchmarks (refer to Appendix -E), indicating its performance advantage. Additionally, our framework shows an average performance improvement of 8.04% and 4.54% on their respective backbones (GCN and SAGE), confirming its general plug-in property. Notably, the sparsification of the adjacency matrix by the NC-mask allows our NCSAGE model to overcome the OOM problem encountered by SAGE on Penn94. All these phenomena highlight the effectiveness of our framework in alleviating heterophily problem, especially the superiority of our proposed NC metric and priors.
@@ VI-B5 RQ5: How do the two hyperparameters of hops and threshold impact NCGCN?
Metric computation and node separation in our framework are controlled by the two key parameters, hops and threshold . To further investigate the impact of these hyperparameters on NCGCN, we fix the other parameters to their optimal settings and plot the test accuracy curves under varying and , as shown in Fig. 9.
When we fix , we observe that NCGCN exhibits relatively small performance fluctuations with varying on the five homophilous graphs and Cora Full. This is because these datasets have relatively stable label distributions, and the NC values computed under different parameter settings are more consistent, resulting in similar node groupings. Consequently, NCGCN is more robust to changes in the hop parameter. In contrast, the neighborhood of different hops in heterophilous graphs will exhibit more diverse label distributions, resulting in significantly distinct node groupings. As a result, NCGCN demonstrates heightened sensitivity to parameter changes. For large-scale heterophilous graphs like Penn94, the 1-hop receptive field already includes a sufficient number of neighbors. The main reason for its poor performance in the 2-hop setting lies in the ineffective utilization of 2-hop information by the target nodes to evaluate their own state.
When we fix , we observe that NCGCN exhibits relatively small performance fluctuations with varying on the five homophilous graphs and Cora Full. This is because these datasets typically have a large proportion of low-NC nodes, as shown in Fig. 5, and as the threshold increases, the node groupings remain stable, leading to consistent model performance. For small-scale heterophilous graphs like Actor, Squirrel Filtered, and Chameleon Filtered, NCGCN is more sensitive to threshold changes due to the more discrete distribution of nodes across different NC values. For larger heterophilous graphs like Penn94, the distribution of nodes with different NC values tends to increase with higher NC values. As the threshold increases, the division between low-NC and high-NC nodes becomes more balanced, which gradually improves the performance of NCGCN ().
@@
ogbn-arxiv | Flickr | |
#Remark | large, temporal | large |
#Classes | 40 | 7 |
#Features | 128 | 500 |
#Nodes | 169343 | 89250 |
#Edges | 1166243 | 334120 |
NH | 0.64 | 0.32 |
NC () | 0.04 | 0.37 |
NC () | 0.06 | 0.57 |
MLP | 55.39 0.17 | 46.91 0.09 |
GCN | 55.87 0.18 | 47.82 0.11 |
SAGE | 57.52 0.29 | 47.87 0.08 |
GAT | 56.81 0.23 | 47.28 0.16 |
H2GCN | 55.23 0.14 | 48.26 0.11 |
GPRGNN | 57.29 0.28 | 48.03 0.10 |
FAGCN | 57.03 0.27 | 48.12 0.36 |
ACM-GCN | 56.78 0.19 | 47.54 0.28 |
CPGNN | OOM | OOM |
WRGAT | OOM | OOM |
GBKSAGE | OOM | OOM |
HOG | OOM | OOM |
GloGNN | OOM | OOM |
SNGNN | 56.21 0.18 | 47.11 0.13 |
CAGNN | 50.34 3.49 | 47.01 0.10 |
NCGCN | 58.29 0.27 | 48.43 0.09 |
NCSAGE | 59.20 0.32 | 48.33 0.13 |
@@ VI-B7 Efficiency Comparison
First, CPGNN, GBKSage and HOG employ two-phase training and soft rewiring techniques, resulting in significant time consumption. WRGAT constructs a weighted multi-relational graph before training, rendering it equivalent to multi-channel GATs with reduced efficiency. These spatial methods typically involve additional computational steps for adjusting message passing rules, such as rewiring, auxiliary tasks, or gating mechanisms, thereby increasing the computational burden. In contrast, the NCGCN architecture is essentially a simple deformation of GCN, while also incorporates the three key designs mentioned in Sec. V-D. Consequently, it avoids excessive computational overhead, ultimately ensuring good computational efficiency. Please refer to Appendix -F for more analysis. Additionally, although some spatial methods consume less computation time, they have certain design flaws in their model architectures. For GloGNN, directly inputting the adjacency matrix into an MLP contradicts permutation invariance. For SNGNN, the approach of selecting the most similar neighbors seems to remain confined within the homophily assumption. CAGNN combines raw and aggregated features without considering label heterogeneity. These design flaws result in these methods falling short in terms of accuracy.
VI-B8 Performance Analysis on Large-scale Graphs with Inductive Learning
In this section, we further analyze the performance of various models on large-scale graphs with inductive learning, focusing on the ogbn-arxiv [56] and Flickr [57] benchmarks. These two datasets present distinct challenges: ogbn-arxiv contains temporal information and a higher degree of homophily, while Flickr is more heterophilous, providing a diverse testing ground for models designed to address the heterophily issue in GNNs. As shown in Table II, the proposed NCGCN and NCSAGE, which are guided by the NC metric, deliver the best performance across both datasets. These results demonstrate the effectiveness of the NC metric in separating nodes based on neighborhood label diversity, allowing the models to apply distinct learning processes to nodes with different levels of neighborhood confusion. This separation is especially crucial for handling the heterophilous nature of Flickr while maintaining high performance on the larger and more homophilous ogbn-arxiv dataset. Additionally, it is important to note that several complex models, such as GBKSAGE and HOG, experience timeout and OOM errors, further demonstrating the challenge of scaling certain spatial methods to large-scale graphs. In contrast, NCGCN and NCSAGE handle these large-scale datasets efficiently, illustrating the scalability of the NC-guided separation framework in real-world graph applications.
@@ VII Conclusions and Outlook
Our framework is simple yet effective enough to be a new start for spatial GNNs. The backbone can be replaced by more message passing models. The robustness and debiasing potential are also open to be explored. Furthermore, NC is an easy-to-compute metric that can effectively separate nodes for different optimization processes. This metric can also be further utilized in data mining, predictability, and interpretability. The limitation of NCGCN is that it utilizes only second-order information following traditional GNNs. It seems sufficient on most real-world datasets except for long-distance dependent ones. Another problem is that high-NC nodes are relatively rare on homophilous graphs, and we can use graph data augmentation [58] for this issue in the future.